# Improving Local Training in Federated Learning via Temperature Scaling

### v2

@@ Abstract
Federated learning is inherently hampered by data heterogeneity: non-i.i.d. training data over local clients. We propose a novel model training approach for federated learning, FLex&\&&Chill, which exploits the Logit Chilling method. Through extensive evaluations, we demonstrate that, in the presence of non-i.i.d. data characteristics inherent in federated learning systems, this approach can expedite model convergence and improve inference accuracy. Quantitatively, from our experiments, we observe up to 6$\times$ improvement in the global federated learning model convergence time, and up to 3.37% improvement in inference accuracy.
@@ 1 Introduction
Federated learning represents a paradigm enabling the training of effective models within distributed environments without explicitly exposing local data McMahan et al. (2017). This approach provides opportunities for communication and computation efficiency during training by sharing model parameters/gradients learned independently from local learning processes while suppressing raw data transmissions. Furthermore, since training operations take place in a decentralized manner, federated learning opens the possibility of parallelizing the model training process via the assistance of local client devices. It presents an attractive solution for handling and exploiting the vast amounts of data generated by many mobile and Internet of Things (IoT) device entities consisting of the overall system Park et al. (2023a).
@@
In contrast to conventional centralized training methods, federated learning faces challenges associated with delayed model convergence and limited performance. A fundamental challenge of federated learning arises from the disparity of training data used at each client device. Federated stochastic optimization encounters difficulties in determining optimal parameters, leading to increased communication overhead between local devices and the central server, as well as the need for additional training on (typically resource-limited) local devices Zhao et al. (2018).
@@
To address these challenges, a number of prior research have focused on optimizing local training operations on federated learning clients or on refining the server-side aggregation process. For instance, FedProx Li et al. (2020) controls the number of local device iterations, aiming to train models resilient to challenges posed by non-independent and non-i.i.d. data environments. SCAFFOLD Karimireddy et al. (2020) achieves expedited convergence and improved model accuracy McMahan et al. (2017) by introducing a correction term in the model aggregation phase to balance the influence of each client. These operations alleviate challenges posed by the non-i.i.d. environment, a common dataset configuration for federated learning clients.
@@
This paper proposes a novel model training approach for federated learning, FLex&\&&Chill (Federated Learning EXploiting logit Chilling), that targets to boost the convergence and model performance via temperature scaling, namely logit chilling, during the local model training operations. Note that a typical neural network generates a probability ($p_{i}$) for each class from its neural network logits ($z_{i}$) by applying a softmax operation (Eq. 1) with the temperature $T$ conventionally set to 1. However, it is not uncommon in other application scenarios to exploit even higher temperature values (i.e., $T\in(1,\infty)$), for example, in knowledge distillation Hinton et al. (2015); Touvron et al. (2021) or neural network calibration Guo et al. (2017).
$\xi(z,i,T)=p_{i}=\frac{exp(z_{i}/T)}{\sum_{j}exp(z_{j}/T)}$ (1)
@@
When employing a temperature $T~{}>$ 1, the probability distribution across classes undergoes smoothing due to the narrowed interval between logits Guo et al. (2017), making such an approach suitable for applications such as knowledge distillation. Notably, despite altering the probabilities, the order of labels’ probability remains unaffected with respect to changes in $T$. While these works successfully explore the impact of varying temperatures for neural network operations, they mainly focus on applying high temperatures after the model training process, mostly for inference operations. Since increasing $T$ reduces the training confidence, such schemes are effective in taming overconfident DNNs.
On the other hand, in federated learning, where disparate training data is common, increasing model confidence of each individually trained client model (using small $T$) prior to parameter aggregation can be a more effective approach. Our investigation involves measuring both the gradient flow and data position shift within the representation space to provide guidelines in this unexplored territory.
@@
Specifically, this work targets to show that utilizing fractional temperature values (i.e., $T\in(0,1)$) during the training process enables effective gradient propagation towards the input layer when applied to federated learning data distribution scenarios. Consequently, we observed that lower temperatures facilitate more efficient data updates within the representation space when dealing with non-i.i.d. datasets, a common characteristic of federated learning systems. From our evaluations with three datasets and three baseline federated learning models, we show that these effects accelerate the federated learning model convergence time by up to $\mathbf{6.00\times}$ and improve the inference accuracy by up to $\mathbf{3.37\%}$, despite potential concerns on model instability when training with with low temperatures. We summarize the contributions of this work in three-fold as the following:
@@
  • To the best of our knowledge, this is the first work that explores the impact of integrating temperature scaling during the training process in a federated learning scenario.
@@
  • We offer empirical evidence that demonstrates the performance enhancements of exploiting logit chilling in federated learning scenarios, by examining the gradient flow and data position shifts within the representation space.
@@
  • We introduce a novel federated learning model training approach, FLex&\&&Chill, showcasing the application of low temperature usage during the training process.
2 Background and Related Work
@@ 2.1 Temperature scaling
In machine learning, a model’s output probability distribution serves diverse purposes. Primarily, classification models offer probabilities for each class, showcasing the model’s confidence in its predictions. This certainty measure allows consideration of the likelihood of each class prior to making a final classification decision. Moreover, this distribution offered by the model helps in understanding the uncertainty of a model on its decisions Park et al. (2023b); Pearce et al. (2021).
@@
For neural networks, computing probabilities from logits often involves employing softmax operations, as in Eq. 1. Here, the temperature parameter ($T>0$) controls the probability distribution’s concentration towards one class, and manages how the categorical probability is dispersed over possible classes Jang et al. (2016). Notably, this categorical probability aligns with the concept of entropy in information theory, offering a perspective where $T$ governs the level of disorder or uncertainty within the distribution. Specifically, with $T\in(1,\infty)$, the probability distribution tends to be smoother as $T$ increases. This arises from the diminished differences between logit values, enhancing the distribution’s entropy. In contrast, $T\in(0,1)$ sharpens the probability distribution, since a lower $T$ amplifies inter-logit differences, effectively polarizing the distribution and concentrating probabilities closer to the maximum logit value. Consequently, model predictions are made more confidently, by allocating higher probabilities to fewer classes, thereby reducing uncertainty in the distribution.
Therefore, temperature scaling operates as a means to adjust the probability distribution within a neural network-based system. Previous work in Knowledge Distillation Hinton et al. (2015); Touvron et al. (2021) exploit the temperature value by applying a high temperature $T\in(1,\infty)$ to achieve a smoothed probability distribution. This manipulation aims to magnify small values within the distribution, often representing hidden information encapsulated within the model. Additionally, in generative models Shih et al. (2023); Wang et al. (2020), temperature scaling regulates the randomness inherent in model outputs. This technique allows for the control of variability levels in generated samples, enabling adjustments to generated output diversity. Temperature scaling is also used in model calibration Guo et al. (2017). Applying temperature adjustments to the model’s output probabilities, typically through a scaling factor $T$, aids in aligning these probabilities more accurately with the true likelihoods or uncertainties present in the data. Likewise, previous work has focused on applying high-temperature values for various purposes, mostly for the purpose of conservatively better understanding the full dataset without bias.
@@ 2.2 Federated learning
In recent years, federated learning has emerged as a promising paradigm in machine learning, particularly in scenarios where data privacy, security, and the decentralization of model training are core concerns. Unlike traditional centralized training approaches, federated learning enables model training across distributed devices or servers while keeping the data localized; thus, addressing privacy challenges associated with centralized data aggregation. This decentralized learning paradigm involves iterative model training where local updates occur on individual devices, and only aggregated model updates are shared with the central server.
Despite its benefits, federated learning poses several challenges. A body of work has tackled the issue of communication overhead stemming from the exchange of frequent model parameter updates. A number of previously proposed works have proposed schemes to effectively abstract the model parameters in various forms to relieve the communication and networking overhead Park et al. (2023a). Furthermore, federated learning systems are prone to facing issues arising from data heterogeneity caused by the diversity or differences in the characteristics, distributions, or formats of data across federated learning clients. Previous works have attempted to address this challenge by addressing relevant issues such as variations in feature/input variations Yu et al. (2021) and spatio-temporal variations Liu et al. (2020).
@@ 3 Design of FLex&Chill
We initiate our discussions by exploring the impact of logit scaling, also known as temperature scaling, providing empirical evidence on the rationale behind our proposed system. Furthermore, in this section, we introduce our system, referred to as FLex&\&&Chill, a novel local training approach designed for efficiently training a federated learning framework.
3.1 Impact of Training with Lower Temperatures
@@
As aforementioned, conventionally, temperature scaling has primarily been focused on calibrating the probability distribution of a model’s output after the model training phase, with a limited exploration into its application during the training process itself. In this section, we present empirical results on three initial experiments specifically crafted to showcase the implications of employing temperature scaling throughout the training phase operations in different dimensions. Through these investigations, our aim is to establish a foundation for comprehending the potential implications and advantages associated with the utilization of temperature scaling techniques in neural network training. We conduct our experiments with the CIFAR10 dataset Krizhevsky et al. (2009) and use a 2-layered CNN architecture McMahan et al. (2017) as the baseline neural network architecture. For experimental details, see APPENDIX B
3.1.1 Gradient flow patterns
@@
Figure 1: Distribution of gradient norm at input layer for correctly/incorrectly inferred samples with varying training temperatures.
@@
Our first experiment focuses on assessing the norm of computed gradients within the input layer for various temperature values ($T$), and we present a sample distribution plot of how the gradients flow for 50K training samples in Figure 2. Note that here we separate the plots for the correctly inferred samples (in the forward pass) and the incorrect ones to observe their differences. The propagation of gradients to the input layer serves as an indicative metric for the model’s learning efficacy. Successful gradient propagation to the input layer implies that the model effectively utilizes and comprehends information from diverse features within the input data during the training phase Evci et al. (2022).
As shown in Figure 2, the mode of distribution shifts to larger values with decreasing temperatures for both correct and incorrect predictions. One reason behind this mode shift stems from the gradient function characteristics of softmax. Eq. 2 presents the gradient of cross-entropy for the softmax activation ($\mathcal{L}$) with varying temperature $T$ and hard target, one-hot encoded labels. For brevity, we present the final derivation result and leave details in the Appendix (see Appendix A). Since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a gradient boost.
$\frac{\partial\mathcal{L}}{\partial{z_{i}}}=\frac{1}{T}(p_{i}-y_{i})=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})$ (2)
3.1.2 Data position shift in the representation space
To better understand the impact of temperature scaling on the training process, we conducted a study into the shift of a data’s position in the representation space (i.e., the distance to the decision boundary). Drawing inspiration from the concept of ‘travel to decision boundary’ as proposed by Kim et al. Kim et al. (2022), we quantified the distance of a data point’s inference result to the decision boundary by determining the magnitude of epsilon ($\epsilon$) value multiplied by the adversarial perturbation Goodfellow et al. (2014) needed to alter the model’s decision.
@@ 3.1.2 Data position shift in the representation space
Figure 2 plots our results in the form of distributions depicting differences between distances to the decision boundary before and after model updates. When higher temperatures ($T={2,4}$) are applied, we can notice that the changes in inference results’ positions are not significant (i.e., plots skewed towards 0). Whereas, with a small $T$, we can notice more pronounced shifts in data positions within the representation space. This suggests that using a small $T$ at the softmax operations can actively steer the model towards (potentially) substantial optimization to better align with the training data.
3.1.3 Visualization in the representation space
To examine the effect of applying a low temperature during the model training phase, we conducted a simple example that classifies 2-dimensional data via logistic regression. We split the data into three different subsets and assign them as three devices with non-i.i.d data. The Softmax layer was applied at different temperatures ($T$=1.0, 0.5) during the local training phase and average model weights for aggregation after 10 local updates.
Figure 3 illustrates the classification boundaries and samples distributed in the representation space for the three clients and the aggregated model. The results suggest that applying $T$ = 0.5 effectively reflects the local dataset characteristics compared to $T$ = 1.0. Specifically, as an example, for clients 0, 1 and the aggregate model, using $T$ = 1.0 fails to make a clear distinction between the red and light green samples; whereas, when applying $T$ = 0.5, this distinction becomes pronounced.
@@ 3.1.3 Visualization in the representation space
Figure 3: Example of data points in the 2D representation space with their respective classification boundaries for different federated learning clients with varying training temperatures.
3.2 FLex&\&&Chill and Logit Chilling
Based on such empirical observations, we hypothesize that training neural networks with lower temperatures can benefit the model convergence and performance for systems exploiting non-i.i.d. data, which is common for federated learning scenarios. In FLex&\&&Chill we term this as logit chilling, as we ‘chill down’ the temperature at the softmax operations, thereby altering the logit computation derived from the neural network during the training process. The focus of a FLex&\&&Chill-based federated learning system is to exploit logit chilling to train a client-side model with low-temperature values. As we observed through the studies above and as our evaluations will show, this has an effect of effectively exploiting crucial information from the training samples to actively alter the model towards the new knowledge.

### v3

@@ Abstract
[…]
keywords:
Federated Learning , Temperature Scaling
journal: Advanced Engineering Informatics
[…]
@@ 1 INTRODUCTION
Federated learning represents a paradigm enabling the training of effective models within distributed environments without explicitly exposing local data [mcmahan2017communication]. This approach provides opportunities for communication and computation efficiency by sharing model parameters and gradients learned independently from local learning processes while suppressing raw data transmissions and potential data leakage [lee2023exploiting, lee2025mind]. Furthermore, since training operations take place in a decentralized manner, federated learning opens the possibility of parallelizing the model training process via the assistance of local client resources. As a result, federated learning presents an attractive solution for handling and exploiting the vast amounts of data generated by many mobile and Internet of Things (IoT) device entities comprising the overall system [park2023attfl, shin2024effective].
@@
However, in contrast to conventional centralized training methods, federated learning faces challenges associated with delayed model convergence and limited performance. More specifically, a fundamental challenge of federated learning arises from the disparity of training data used at each client device. Given this inevitable training data disparity rising from its distributed nature, federated stochastic optimization encounters difficulties in determining optimal parameters, which can lead to increased communication overhead between local devices and the central server, as well as needing additional training on (typically resource-limited) local devices [zhao2018federated].
@@
To address this challenge, several prior work have proposed schemes to optimize local training operations for federated learning clients or have targeted to refine the server-side model aggregation operations. For instance, FedProx [li2020federated] appends a proximal term, aiming to train models resilient to challenges posed by non-independent and non-i.i.d. data environments. SCAFFOLD [karimireddy2020scaffold] achieves expedited convergence and improved model accuracy by introducing a correction term in the local model training phase to balance the influence of each client. These operations alleviate challenges posed by the non-i.i.d. data environment, a common data distribution for federated learning clients.
@@
We note that a potentially orthogonal approach to these prior efforts is to exploit the fact that a typical neural network generates a probability ($p_{i}$) for each target class from its neural network logits ($z_{i}$) by applying a softmax operation (Eq. 1) with a temperature $T$ value conventionally set to 1. However, it is not uncommon to exploit even higher temperature values (i.e., $T\in(1,\infty)$) in use cases such as knowledge distillation [hinton2015distilling, touvron2021training] or neural network calibration [guo2017calibration].
$p_{i}=\frac{exp(z_{i}/T)}{\sum_{j}exp(z_{j}/T)}$ (1)
@@
When employing a temperature $T>1$ for scaling the logits, the probability distribution across classes undergoes smoothing due to the narrowed inter-logit intervals [guo2017calibration], making higher temperature adjustments suitable for applications such as knowledge distillation. This is possible since, despite altering the model’s internal probabilities, the order of labels’ probabilities remains unaffected with respect to changes in $T$.
While the use of diverse temperatures for neural network operations has been shown to be effective, works presented until now mainly focus on applying high temperatures after the model is trained, making it mostly applicable for the inference operations only. For instance, since increasing $T$ reduces the model confidence, such schemes are effective in curbing the overconfidence of DNNs [guo2017calibration]. On the other hand, the effect of exploiting temperature scaling during the training process has gathered relatively limited attention in the current literature [agarwala2020temperature]. Despite the distinctive characteristics of federated learning scenarios (e.g., data disparity and training parallelism), the benefits of employing temperature scaling during the federated training operations are left understudied.
This paper proposes a novel model training approach for federated learning, FLex&\&Chill (Federated Learning EXploiting logit Chilling), that targets to boost the model convergence and performance via temperature scaling, namely logit chilling, during the local model training operations. This paper analyzes the efficacy and effectiveness of employing temperature scaling during the local training process in a federated learning scenario via theoretical convergence analysis and empirical evidence to provide guidelines in this unexplored territory.
@@
Specifically, we target to show that utilizing fractional temperature values (i.e., $T\in(0,1)$) during model training enables effective gradient propagation towards the input layer when applied to federated learning with non-i.i.d. datasets. From our evaluations with three datasets and three baseline federated learning models, we show that these effects accelerate the federated learning model convergence time by up to $\mathbf{6.00\times}$ and improve the inference accuracy by up to $\mathbf{3.37\%}$, despite potential concerns on model instability when training with low temperatures. Specifically, we summarize the contributions of this work in three folds as follows:
@@
  • To the best of our knowledge, this is the first work that assesses the impact of integrating temperature scaling during the training process in a federated learning scenario.
@@
  • We offer theoretical explanations and empirical evidence that demonstrate the performance enhancements of exploiting logit chilling in federated learning scenarios.
@@
  • We introduce a novel federated learning model training approach, FLex&\&Chill, showcasing the application of low-temperature usage during the training process.
The remainder of this work is structured as follows. The following section will present related work in this domain to position our work among previous work. Section 3 presents the design of FLex&\&Chill along with deep analysis of its operations. We will present experimental results using three datasets in Section 4 and suggest interesting discussion points in Section 5. Finally, we conclude the paper in Section 6. We note that an open source implementation of the proposed scheme can be accessed at https://github.com/eis-lab/temperature-scaling.
2 RELATED WORK
@@ 2.1 Temperature Scaling
In machine learning, a model’s output probability distribution serves diverse purposes. Primarily, classification models offer a probability for each class, showcasing the model’s confidence in its predictions. This model-generated certainty measure allows consideration of the likelihood of each class prior to making a final classification decision. Moreover, this distribution offered by the model helps in understanding the uncertainty of a model on its decisions [park2023self, pearce2021understanding].
Figure 1: Effect of varying temperature $T$ on the output distribution of the softmax function, illustrating how lower $T$ sharpens class probabilities and higher $T$ produces smoother, more uniform distributions. Best viewed in color.
@@
For neural networks, computing label probabilities from logits often involves softmax operations, which is a special case ($T=1$) of a Boltzmann distribution as in Eq. 1. Here, the temperature parameter ($T>0$) controls the concentration of probability towards the class with the largest logit and manages how the categorical probability is dispersed over possible classes [jang2016categorical]. Figure 1 shows how the per-class probability distribution changes for varying temperature (i.e. $T$) values. Specifically, with $T\in(1,\infty)$, the probability distribution tends to be smoother as $T$ increases (green line in Fig. 1). This arises from the diminished ratio between exponentials of equally scaled values. In contrast, $T\in(0,1)$ shifts the probability distribution into a point-mass, as a lower $T$ amplifies the ratio of the exponential of scaled logits, effectively polarizing the distribution and concentrating probabilities to the maximum logit value (blue line in Fig. 1). Consequently, model predictions are made more confidently by allocating higher probabilities to fewer classes. Therefore, temperature scaling operates as a means to adjust the probability distribution within a neural network-based system.
Temperature scaling was applied in previous works for knowledge distillation [hinton2015distilling, touvron2021training], generative models [shih2023long, wang2020contextual], gradient analysis [lee2024detrigger] and model calibration [guo2017calibration]. Applying temperature adjustments to the model’s output probabilities, typically through a scaling factor $T$, aids in aligning the probabilities more accurately with the true likelihoods or uncertainties present in the data.
However, most prior works have focused primarily on applying temperature scaling as a post-training technique for probability calibration, leaving its effects during training largely unexplored. While pioneering research [agarwala2020temperature] investigated temperature scaling during training and showed its positive impact in model generalization, their analysis was limited to a centralized learning scenario. Furthermore, works such as FedRS propose changing local objective functions to apply temperature scaling in federated model training operations [li2021fedrs]. However, the work is rather heuristic and does not present a thorough theoretical explanation of why altering the softmax function can improve convergence and mitigate data heterogeneity. In contrast, our work explicitly highlights (i) the effect of temperature scaling applied during the training phase rather than post-training, and (ii) its impact specifically in federated learning scenarios characterized by heterogeneous, non-i.i.d. data distributions.
@@ 2.2 Federated Learning
Recently, federated learning has emerged as a promising paradigm in machine learning, particularly in scenarios where data privacy, security, and the decentralization of model training are core concerns. Unlike traditional centralized training approaches, federated learning enables model training across distributed devices or servers while keeping the data localized; thus, addressing privacy challenges associated with centralized data aggregation. This decentralized learning paradigm involves iterative model training where local updates occur on individual devices, and only aggregated model updates are shared with the central server. Specifically, federated learning aims to solve the following optimization problem:
$\min_{w}F(w)=\sum_{k=1}^{K}p_{k}F_{k}(w),$
where $K$ denotes the total number of clients participating in training, $w$ represents the global model parameters to be optimized, $p_{k}\geq 0$ is the weight corresponding to the $k$-th client (typically proportional to the size of the local dataset), and $F_{k}(w)$ denotes the local objective function defined as follows:
$F_{k}(w)=\frac{1}{|D_{k}|}\sum_{(x,y)\in D_{k}}\ell(w;x,y),$
where $D_{k}$ is the local dataset on client $k$, and $\ell(w;x,y)$ represents the loss function evaluating the prediction of the model $w$ for data sample $(x,y)$. The key challenge of federated learning lies in optimizing the global objective $F(w)$ under the constraint that the clients’ data remains distributed and private, each client trains locally without directly sharing its raw data.
Despite its benefits and active research, federated learning poses several challenges. A body of work has tackled the issue of communication overhead stemming from the exchange of frequent model parameter updates. A number of previously proposed works have proposed schemes to effectively abstract the model parameters in various forms to relieve the communication and networking overhead [park2023attfl]. Furthermore, federated learning systems are prone to facing issues arising from data heterogeneity caused by the diversity or differences in the characteristics, distributions, or formats of data spread across distributed federated learning clients. Previous works have attempted to address this challenge by addressing relevant issues such as variations in feature/input variations [yu2021fed2] and spatio-temporal variations [liu2020client].
@@ 3 DESIGN OF FLEX&CHILL
We begin discussing our framework design by examining the effect of temperature scaling (also referred to as logit scaling), providing both theoretical and empirical justification for our approach. On the theoretical side, we analyze how temperature scaling influences training dynamics in federated learning, with a particular focus on gradient amplification and its implications for convergence. Empirically, we investigate how applying temperature scaling alters the data distribution and decision boundaries within the representation space. Building on these insights, we introduce FLex&\&Chill, a novel local training strategy tailored to enhance the efficiency and performance of federated learning systems.
3.1 Theoretical Analysis
@@
As aforementioned, conventionally, temperature scaling has primarily been focused on calibrating the probability distribution of a model’s output after the model training phase, with limited exploration into its application during the training process itself. In this section, we provide a theoretical analysis of the effect of temperature scaling during training with respect to the convergence in a federated learning scenario. Based on this investigation, we aim to establish a foundation for comprehending the potential implications and advantages associated with the utilization of temperature scaling techniques in neural network training.
$\frac{\partial\mathcal{L}}{\partial{z_{i}}}=\frac{1}{T}(p_{i}-y_{i})=\frac{1}{T}\left(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i}\right)$ (2)
Eq. 2 presents the gradient of cross-entropy for the softmax activation ($\mathcal{L}$) with varying temperature $T$ and hard target, one-hot encoded labels. For brevity, we present the final derivation result and present detailed proof in the Appendix. Since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a boost in the magnitude of the gradient. Furthermore, from the perspective of the loss term $(p_{i}-y_{i})$, when the model yields an incorrect decision, the norm of the loss term significantly increases as $p_{i}\xrightarrow{}0$.
Building upon the theoretical analysis on the convergence of federated learning by Li et al. [Li2020On], we can derive the inequality Eq. 3, where $F(w^{t})$ is the loss of the neural network with the weight ($w^{t}$) at round $t$, Lipschitz constant $L$, the convexity constant $\mu$, and the upper bound of variance of the stochastic gradient at local devices $\sigma^{2}$. The inequality suggests that the difference between the optimal weight $w^{*}$ and the actual weight ($w_{t}$) can be minimized by iteratively updating the weight.
$\begin{split}&\mathbb{E}[F(w^{(t+1)})-F(w^{*})]\leq\\ &(1-\frac{\eta\mu}{T})\mathbb{E}[F(w^{(t)})-F(w^{*})]+\frac{L\eta^{2}\sigma^{2}}{2T^{2}}.\end{split}$ (3)
Here, we can identify the role of softmax temperature ($T$) regarding the convergence of federated settings. In the inequality, the temperature $T$ controls the speed of convergence. For instance, a high temperature ($T>1$) slows the convergence as it reduces the term $\frac{\eta\mu}{T}$. Furthermore, $T$ also affects the stability of convergence as it controls the variance term ($\frac{L\eta^{2}\sigma^{2}}{2T^{2}}$). $T>1$ also reduces the amplitude of the variance term, which consequently stabilizes the convergence (Proof in Appendix). Based on these findings, we can consider the temperature as a key hyperparameter in federated learning. We note that the above discussion can be extended to more complex adaptive optimization methods. For brevity, these details are provided in the Appendix.
Besides the mathematical analysis, to provide an intuitive rationale for applying temperature scaling during local training, the low temperatures effectively penalize incorrect predictions strongly while rewarding correct ones, enabling aggressive learning focused on challenging samples. This intuition aligns with the experimental results presented in the following section. However, it is important to note that excessively aggressive updates may cause instability.
3.2 Temperature v.s. Learning Rate
Since $\eta$ and $T$ have a reciprocal relationship in Eq. 3, we can question the relationship between the learning rate ($\eta$) and the temperature ($T$). Specifically, Will decreasing or increasing the temperature be equivalent to increasing or decreasing the learning rate? We emphasize that while temperature and learning rate exhibit similar effects on convergence, they are not identical.
$w_{t+1}$ $=w_{t}-\eta\nabla f(w_{t})$ (4)
$=w_{t}-\frac{\eta}{T}\left(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i}\right)$ (5)
Eq. 4 presents the stochastic update of a neural network parameter $w_{t}$ at step $t$ with learning rate $\eta$. In Eq. 5, the update rule can be written by introducing the softmax temperature parameter $T$. Dividing the learning rate $\eta$ by the temperature $T$ suggests that increasing $T$ would have a similar effect to reducing $\eta$, and vice versa. $T$ has, however, an additional effect in calculating the Boltzmann distribution inside the parentheses.
$\eta$ only controls the magnitude of the parameter update, affecting the overall convergence dynamics. On the other hand, $T$ has two-fold effects: changing convergence dynamics and modulating the output distribution’s sensitivity. $T$ impacts the relative importance of individual gradients and their direction rather than only uniformly scaling the update size. Therefore, although $\eta$ and $T$ appear to play interchangeable roles for some cases, the underlying mechanisms through which they influence the training process are fundamentally different.
3.3 Empirical Analysis
In this section, we present empirical results via three initial experiments specifically crafted to support the analytical results and showcase the implications of employing temperature scaling throughout the training phase operations in different dimensions. Through these investigations, we aim to establish a foundation for comprehending the potential implications and advantages associated with the utilization of temperature scaling techniques in neural network training. We conduct our experiments with the CIFAR10 dataset [krizhevsky2009learning] and use a 2-layered CNN architecture [mcmahan2017communication] as the baseline neural network architecture.
@@ 3.3 Empirical Analysis
Figure 2: Distribution of gradient norm at input layer for correctly (top) / incorrectly (bottom) inferred samples with varying training temperatures.
Input gradient norm. Our first experiment focuses on assessing the norm of computed gradients within the input layer for various temperature values ($T$), and we present a sample distribution plot of the norm of input layer gradients for 50K training samples in Figure 2. Note that here we separate the plots for the correctly inferred samples (in the forward pass) and the incorrect ones to observe their differences. The propagation of gradients to the input layer serves as an indicative metric for the model’s learning efficacy. Successful gradient propagation to the input layer implies that the model effectively utilizes and comprehends information from diverse features within the input data during the training phase [evci2022gradient].
As shown in Figure 2, the mode of distribution shifts to larger values with decreasing temperatures for both correct and incorrect predictions. As mentioned, the reason behind this mode shift stems from the gradient function characteristics of softmax. As shown in Eq. 2, since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a gradient boost.
@@
Data position shift in the representation space. To better understand the impact of temperature scaling on the training process, we conducted a study into the shift of a data’s position in the representation space (i.e., the distance to the decision boundary). Drawing inspiration from the concept of travel to decision boundary as proposed by Kim et al. [kim2022curved], we quantified the distance of a data point’s inference result to the decision boundary by determining the magnitude of epsilon ($\epsilon$) value multiplied by the adversarial perturbation [goodfellow2014explaining] needed to alter the model’s decision.
@@
Figure 3 plots our results in the form of distributions depicting differences between distances to the decision boundary before and after model updates. When higher temperatures (T=2,4T={2,4}) are applied, we can notice that the changes in inference results’ positions are not significant (i.e., plots skewed towards 0). Whereas, with a small $T$, we can notice more pronounced shifts in data positions within the representation space. This suggests that using a small $T$ at the softmax operations can actively steer the model towards (potentially) substantial optimization to better align with the training data.
Visualization in the representation space
@@
Figure 4: Example of data points in the 2D representation space with their respective classification boundaries for different federated learning clients with varying training temperatures. Best viewed in color.
To examine the effect of applying a low temperature during the model training phase, we conducted a simple example that classifies 2-dimensional data via logistic regression. We split the data into three different subsets and assigned them as three devices with non-i.i.d. data. The Softmax layer was applied at different temperatures (TT=1.0, 0.5) during the local training phase and average model weights for aggregation after 10 local updates.
Figure 4 illustrates the classification boundaries and samples distributed in the representation space for the three clients and the aggregated model. The results suggest that applying $T$ = 0.5 effectively reflects the local dataset characteristics compared to $T$ = 1.0. Specifically, as an example, for clients 0, 1, and the aggregate model, using $T$ = 1.0 fails to make a clear distinction between the red and light green samples; whereas, when applying $T$ = 0.5, this distinction becomes pronounced.
3.4 FLex&\&Chill and Logit Chilling
Based on the theoretical analysis and our empirical observations, we find that training neural networks with lower temperatures directly amplifies the gradient magnitude and accelerates convergence in the presence of non-i.i.d. data, which is a typical regime for federated learning. In particular, our analysis shows that the softmax cross-entropy gradient scales with the factor $1/T$, and the resulting convergence bound includes the term $\eta\mu/T$, indicating faster convergence when $T<1$. Complementing this, our empirical studies in the representation space (e.g., Figures 3 and 4) show that lower temperatures move samples farther from the decision boundary, especially for the misclassified points, leading to more aggressive and influential local updates at each client.
We formulate these insights as an explicit optimization strategy, termed FLex&\&Chill. Specifically, FLex&\&Chill exploits logit chilling by “chilling down” the temperature at the softmax layer during local training on each client, thereby sharpening the output distribution and amplifying gradients throughout the network. The focus of a FLex&\&Chill-based federated learning system is to leverage these low-temperature updates so that each client can more strongly encode new knowledge into its local model, and the subsequent aggregation of these models at the server accelerates the global convergence toward a high-quality solution. As our evaluations in the following sections demonstrate that FLex&\&Chill consistently achieves faster convergence and improved accuracy compared to conventional training with $T=1$ under various non-i.i.d. federated learning scenarios.
We formulate the operations of FLex&\&Chill using Algorithm 1. Here, typical federated learning operations take place at both the server and clients except for the fact that the server determines a target $T$ at each round and local model training operations are performed with $T$. Since FLex&\&Chill only manipulates the local training process which inherently provides orthogonality to our proposed method, it is extremely easy to integrate with existing federated learning frameworks (e.g., FedProx [li2020federated], SCAFFOLD [karimireddy2020scaffold], FedBN [li2021fedbn]).
Algorithm 1 FLex&\&Chill
Data ($D_{1},D_{2},...,D_{N}$) where $D_{i}$ is the user $i$’s local data.
Model ($w^{1}_{t},w^{2}_{t},...,w^{N}_{t}$) where $w_{t}^{i}$ is the user $i$’s model in $t$-th round.
Parameter: Temperature $T$
  Server executes:
  Initialize $w_{0}$
  for each round t=1,2,,Rt=1,2,...,R do
   $S_{t}\xleftarrow{}$ (random set of $m$ clients)
   for $k\in S_{t}$ do
    $w_{t}^{k}\xleftarrow{}$ ClientUpdate($w_{t},T$)
   end for
   $w_{t+1}\xleftarrow{}$ Aggregation($S_{t}$)
  end for
  ClientUpdate($w_{t}$, $T$):
  $w_{t}^{k}\xleftarrow{}w_{t}$
  for data, label $\in D_{k}$ do
   $\mathbf{z}=w_{t}^{k}(data)/T$
   $l(\mathbf{z})=loss(\mathbf{z};label)$
   $w_{t}^{k}\xleftarrow{}LocalUpdate(w_{t}^{k},l(\mathbf{z}))$
  end for
  Return $w_{t}^{k}$ to server